# Using Frame Randomization to Mitigate Errors in Quantum Optimization

### v1

@@ Abstract
Error mitigation is essential for near-term quantum devices, and one promising technique is frame randomization. This method inserts random twirling gates into a circuit to reduce errors while preserving unitarity and depth. We apply frame randomization to the quantum approximate optimization algorithm (QAOA) with $p=1$ on a superconducting quantum circuit system, demonstrating its potential to improve energy calculations. Specifically, we investigate the use of QAOA to calculate the lowest energy state of a frustrated Ising ring system and compare the results of randomized circuits generated using two different randomized techniques. Our results show that both methods can mitigate errors, with expected extremal energy values of $5.25\pm 0.145$ and $4.08\pm 0.36$, for Randomized Compilation and Pauli frame randomization respectively, compared to $2.63\pm 0.068$ without randomization and $5.676\pm 0.006$ with a noiseless simulator.
@@ I Introduction
Quantum computing has the potential to revolutionize various fields, but it is hindered by the presence of noise in quantum systems. To mitigate this noise, techniques such as quantum error correction, suppression, and mitigation are employed [1] [2] [3]. Quantum error correction is impractical for current noisy intermediate-scale quantum (NISQ) devices due to high error rates [4]. Therefore, in this paper we focus on an error mitigation strategy, frame randomization, which has been shown to reduce errors in quantum computations [5][6]. In this paper, we explore the application of frame randomization to the Quantum Approximate Optimization Algorithm (QAOA) for solving a frustrated Ising ring problem.
I.1 Frame Randomization
Pauli frame randomization and Randomized Compilation are very similar error mitigation techniques that mitigate errors associated with operating two-qubit gates [6]. Here we will refer to the general approach that encompasses both techniques as frame randomization.
To apply frame randomization to a specific circuit, we start by drawing the circuit diagram and framing different cycles within the circuit. For example, in the circuit in Figure 1, we define 4 cycles and each cycle in this example contains no more than one gate operation for each qubit.
@@ I.1 Frame Randomization
Two slight variations of the general technique of frame randomization are Pauli frame randomization, which inserts random Pauli gates, and Randomized Compilation, which inserts random Clifford gates. [6]. We will investigate Pauli Frame Randomization by using the TKET software package and Randomized Compilation by using the True-Q software package.
@@ II Experimental Setup
To compare errors, we run these circuits on a closed/ideal classical simulator, a noisy simulator that used a device-specific noise model, and on quantum processors. We then apply Pauli frame randomization and Randomized Compilation and run the modified circuits on quantum processors.
@@ II.1 Randomized Compilation with True-Q
Using the True-Q package, we implement our circuit and define it to contain two randomized cycles because the entangling RZZ gates were implemented in two cycles. We generate randomized circuits and run them on IBM quantum computers. The True-Q package chooses randomized gates from the group of Clifford gates [6].
II.2 Pauli Frame Randomization with TKET
We also use TKET to implement Pauli frame randomization. In the same way, we build our circuit and define it to contain two cycles. We generate randomized circuits and run them on IBM quantum computers. The TKET package chooses randomized gates from the group of Pauli gates [7]. A circuit generated using Pauli frame randomization is shown in Figure 10 in the Appendix.
@@ III Results
Figure 9: Energy Landscape generated by running QAOA on ibm_hanoi with 10 compilations using TKET Pauli frame randomization (Spring 2024).
Randomized Compilation and Pauli Frame Randomization
We apply frame randomization to the same collection of circuits as before and compare the energy landscapes generated. We choose to apply 20 compilations because this was expected to provide a reasonable improvement in results without too much additional computation [6]. Results generated using True-Q are plotted in Figure 8. Results generated using TKET are plotted in Figure 9.
@@ IV Discussion
We use QAOA to solve for the lowest energy state of a 12-node frustrated Ising ring system, implementing it on a classical quantum simulator and IBM’s superconducting quantum computers. On IBM’s devices, we run these circuits with no error mitigation, using True-Q Randomized Compilation, and using TKET Pauli frame randomization. Across all runs, the quantum computer produces the same periodic pattern of high and low energy values in the energy landscape, although the simulator achieves a larger range of expected energy values due to its lack of noise. We observe a significant improvement in expected energy values when applying True-Q Randomized Compilation and TKET Pauli frame randomization, as shown in Figures 5-9.
We conclude that frame randomization shows promise for minimizing errors in QAOA circuits, producing energy expectation values closer to those of the noiseless circuit than unmitigated circuits.
In the future, we plan to explore the combined effects of additional optimization techniques and frame randomization on the resulting energy values. We also intend to investigate how different numbers of random compilations impact the output.
@@ Acknowledgements.
This paper is adapted from an undergraduate thesis submitted to the Glynn Family Honors Program at the University of Notre Dame. We acknowledge the contributions of Anthony Hoffman, advisor for Rachel’s thesis at Notre Dame, Jarrett Smalley, initial code developer, and Arnaud Carignan-Dugas of Keysight Technologies for assistance with the True-Q software. We also thank Lockheed Martin for funding this project.

### v2

# Using Universal Frame Randomization and Randomized Compilation to Mitigate Errors in Quantum Optimization

@@ Abstract
Error mitigation is essential for near-term quantum devices, and two promising techniques are universal frame randomization and Randomized Compilation. These methods insert random twirling gates into a circuit to reduce errors while preserving unitarity and depth. We apply universal frame randomization and Randomized Compilation to the quantum approximate optimization algorithm (QAOA) with $p=1$ on a superconducting quantum circuit system, demonstrating its potential to improve energy calculations. Specifically, we investigate the use of QAOA to calculate the lowest energy state of a frustrated Ising ring system and compare the results of randomized circuits generated using both techniques. Our results show that both methods can mitigate errors, with expected extremal energy values of $5.25\pm 0.145$ and $4.08\pm 0.36$, for Randomized Compilation and universal frame randomization respectively, compared to $2.63\pm 0.068$ without randomization and $5.676\pm 0.006$ with a noiseless simulator.
@@ I Introduction
Quantum computing has the potential to revolutionize various fields, but it is hindered by the presence of noise in quantum systems. To mitigate this noise, techniques such as quantum error correction, suppression, and mitigation are employed [1] [2] [3]. Quantum error correction is impractical for current noisy intermediate-scale quantum (NISQ) devices due to high error rates [4]. Therefore, in this paper we focus on error mitigation strategies, universal frame randomization and Randomized Compilation, which have been shown to reduce errors in quantum computations [5][6]. In this paper, we explore the application of universal frame randomization and Randomized Compilation to the Quantum Approximate Optimization Algorithm (QAOA) for solving a frustrated Ising ring problem.
I.1 Universal Frame Randomization & Randomized Compilation
Universal frame randomization and Randomized Compilation are similar error mitigation techniques that mitigate errors associated with operating two-qubit gates [6].
To apply either of these techniques to a specific circuit, we start by drawing the circuit diagram and framing different cycles within the circuit. For example, in the circuit in Figure 1, we define 4 cycles and each cycle in this example contains no more than one gate operation for each qubit.
@@ I.1 Universal Frame Randomization & Randomized Compilation
Both universal frame randomization and Randomized Compilation can mitigate errors for circuits with a universal gate set. [6]. Universal frame randomization allows cycles to contain any combination of single and two-qubit gates, while Randomized Compilation distinguishes between these types of gates and does not allow two qubit gates inside cycles.
We will investigate universal frame randomization by using the TKET software package and Randomized Compilation by using the True-Q software package.
@@ II Experimental Setup
To compare errors, we run these circuits on a closed/ideal classical simulator, a noisy simulator that used a device-specific noise model, and on quantum processors. We then apply universal frame randomization and Randomized Compilation and run the modified circuits on quantum processors.
@@ II.1 Randomized Compilation with True-Q
Using the True-Q package, we implement our circuit and define it to contain two randomized cycles because the entangling RZZ gates were implemented in two cycles. We generate randomized circuits and run them on IBM quantum computers.
II.2 Universal Frame Randomization with TKET
We also use TKET to implement universal frame randomization. In the same way, we build our circuit and define it to contain two cycles. We generate randomized circuits and run them on IBM quantum computers. A circuit generated using universal frame randomization is shown in Figure 10 in the Appendix.
@@ III Results
Figure 9: Energy Landscape generated by running QAOA on ibm_hanoi with 10 compilations using TKET universal frame randomization (Spring 2024).
Randomized Compilation and Universal Frame Randomization
We apply Randomized Compilation and universal frame randomization to the same collection of circuits as before and compare the energy landscapes generated. We choose to apply 20 compilations because this was expected to provide a reasonable improvement in results without too much additional computation [6]. Results generated using True-Q are plotted in Figure 8. Results generated using TKET are plotted in Figure 9.
@@ IV Discussion
We use QAOA to solve for the lowest energy state of a 12-node frustrated Ising ring system, implementing it on a classical quantum simulator and IBM’s superconducting quantum computers. On IBM’s devices, we run these circuits with no error mitigation, using True-Q Randomized Compilation, and using TKET universal frame randomization. Across all runs, the quantum computer produces the same periodic pattern of high and low energy values in the energy landscape, although the simulator achieves a larger range of expected energy values due to its lack of noise. We observe a significant improvement in expected energy values when applying True-Q Randomized Compilation and TKET universal frame randomization, as shown in Figures 5-9.
We conclude that universal frame randomization and Randomized Compilation show promise for minimizing errors in QAOA circuits, producing energy expectation values closer to those of the noiseless circuit than unmitigated circuits.
In the future, we plan to explore the combined effects of additional optimization techniques with universal frame randomization or Randomized Compilation on the resulting energy values. We also intend to investigate how different numbers of random compilations impact the output.
@@ Acknowledgements.
This paper is adapted from an undergraduate thesis submitted to the Glynn Family Honors Program at the University of Notre Dame. We acknowledge the contributions of Anthony Hoffman, advisor for Rachel’s thesis at Notre Dame, Jarrett Smalley, initial code developer, Arnaud Carignan-Dugas of Keysight Technologies for assistance with the True-Q software, and Joseph Emerson for feedback on the first version of this paper. We also thank Lockheed Martin for funding this project.